OpenStudy (anonymous):

Find the perimeter and the area of a right triangle if the shortest side is 5cm and the longest side is 13cm

OpenStudy (ash2326):

@tthomas0557 do you know the Pythagoras theorem?

OpenStudy (anonymous):

No I sure don't

OpenStudy (ash2326):

OK, I'll explain you :)

OpenStudy (anonymous):

Thank You :)

OpenStudy (ash2326):

Pythagoras theorem states that for a right triangle. the square of the hypotenuse ( largest side) is equal to the sum of the squares of the other two sides. \[AC^2=AB^2+BC^2\]  Do you understand this?

OpenStudy (anonymous):

Yes I am following!!

OpenStudy (ash2326):

Find the perimeter and the area of a right triangle if the shortest side is 5cm and the longest side is 13cm Which is the longest side here?

OpenStudy (anonymous):

AC?

OpenStudy (ash2326):

good, now the shortest side could be either AB or BC, it doesn't matter here. Let BC=5 Can you find AB?

OpenStudy (anonymous):

Forgive me but see... I was following right until you asked me if I coulf find AB.. Its like I don't know what I should do....

OpenStudy (ash2326):

Pythagoras theorem states that for a right triangle. the square of the hypotenuse ( largest side) is equal to the sum of the squares of the other two sides. \[AC^2=AB^2+BC^2\]

OpenStudy (anonymous):

Oh Ok... can you give me a sec. I think I know let me work it out and I will give you an answer....

OpenStudy (anonymous):

would it happen to be 16?

OpenStudy (ash2326):

check again :)

OpenStudy (anonymous):

hummm ok... let me see

OpenStudy (anonymous):

Ok.... I tried a different way of working out the problem... what about 144?

OpenStudy (ash2326):

\[AC^2=144\] so \[AC=\sqrt{144}\]

OpenStudy (anonymous):

I have a super silly question but that symbol means?? I have not done this type of math is several years. I am a little rusty!

OpenStudy (anonymous):

It just means that your anser has been squared correct?

OpenStudy (ash2326):

square root of 144

OpenStudy (anonymous):

But the answer is 144 correct?

OpenStudy (ash2326):

square root of 144= 12 cause \[12^2=144\]

OpenStudy (anonymous):

ohhhhh So the answer is 12?? I forgot to square the 144

OpenStudy (ash2326):

yeah AB=12 can you find the are and perimeter now?

OpenStudy (anonymous):

ewww. Ok.. I think I can figure it out.. Lets see..

OpenStudy (ash2326):

cool :) I'm here to help you in case you need my help

OpenStudy (anonymous):

ok.. SO I have that 30 is the area and also for the perimeter???

OpenStudy (anonymous):

@ash2326 ok.. SO I have that 30 is the area and also for the perimeter???

OpenStudy (ash2326):

yeah, you're correct:D